Does { a + b √ 2 | a, b ∈ Z } form a ring (with the usual operations in R )?


Is it commutative? Does it have a unity? Be sure to justify your answers.



1
Expert's answer
2021-09-06T16:06:04-0400

"\text{The set $A = \{ a + b\sqrt2\}$ is a subring of $R$ since for any arbitrary elements: }\\
a + b\sqrt2, c+ d\sqrt2\\
a + b\sqrt2 - c+ d\sqrt2 = (a-c) + (b-d)\sqrt2 \in A \\
\text{and } (a + b\sqrt2) (c+ d\sqrt2) = (ac + 2bd)(bc + ad)\sqrt2 \in A \\
\text{Hence since $A$ form a subring, thus it suffices to say that $A$ form a ring.} \\
\text{$A$ is commutative since:} (a + b\sqrt2) (c+ d\sqrt2) = (c+ d\sqrt2)(a + b\sqrt2) = (ac + 2bd)(bc + ad)\sqrt2 \in A \\
\text{It has a unity since $1 \in A$ and } 1\cdot(a + b\sqrt2) = (a + b\sqrt2) \cdot 1= (a + b\sqrt2)"